OpenStudy (anonymous):

2(4x - 3) - 8 = 4 + 2x

OpenStudy (anonymous):

1/3 = x? right?

jimthompson5910 (jim_thompson5910):

2(4x-3)-8=4+2x 8x-6-8=4+2x 8x-14=4+2x 8x=4+2x+14 8x-2x=4+14 6x=4+14 6x=18 x=(18)/(6) x=3

OpenStudy (anonymous):

Here are the steps in order :D... 2(4x-3)-8=4+2x After distributing... 8x-6-8=4+2x Then subtract like terms... 8x-14=4+2x Now simplify the equation... 6x=18 x=3